OpenStudy (anonymous):

Point T is the midpoint of JH. The coordinate of T is (0, 5) and the coordinate of J is (0, 2). The coordinate of H is: (0, 8) (0, 3) (0, 7) (0, 11)

OpenStudy (anonymous):

if (a,b)= midpoint of JH which is T, THEN (a,b)= { (c+0) /2 , ( d+2) /2 } however (a,b) is given and is ( 0,5) so 0= c+0 /2 and 5= ( d+2 ) / 2 , solve for c and solve for d